A river has a current of 4mph. Find the speed of Simon's boat in still water if it goes 40 miles downstream in the same time as 24 miles upstream.
Answer by checkley77(12569)

D=RT
40=(X+4)T OR T=40/(X+4)
24=(X-4)T OR T=24/(X-4)
40/(X+4)=24/(X-4) CROSS MULTIPLY.
40(X-4)=24(X+4)
40X-160=24X+96
40X-24X=96+160
16X=256
X=256/16
X=16 MPH IS THE SPEED OF THE BOAT.
PROOF:
40/(16+4)=24/(16-4)
40/20=24/12
2=2